OpenStudy (anonymous):

Find the limit of the function algebraically. lim(x->-3) x^2-9/x^3+3

OpenStudy (anonymous):

-18/-24

OpenStudy (anonymous):

is it really \[\frac{x^2-9}{x^3+3}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if so the limit is just zero the numerator is zero and the denominator is not

OpenStudy (anonymous):

\[\frac{(-3)^2-9}{(-3)^2+3}=\frac{0}{-24}\]

OpenStudy (anonymous):

omg my calculator keeps giving me -18

OpenStudy (anonymous):

otherwise known as \(0\)

OpenStudy (anonymous):

that is because your calculator cannot read your mind

OpenStudy (anonymous):

if you type \[-3^2\] in your calculator you will get \(-9\)

OpenStudy (anonymous):

okay my bad

OpenStudy (anonymous):

but it should be \((-3)^2=9\)